# Basic representations of genus zero nonabelian Hodge spaces

### v2

@@ 4.4. Degenerate Painlevé III
4.5. Doubly degenerate Painlevé III
The Lax representation for the doubly degenerate Painlevé equation has two irregular singularities, one of them at infinity, both corresponding to a second order pole, but with a nilpotent leading term at each pole so that at each pole, in the Turritin-Levelt decomposition, there is a single twisted Stokes circle of slope 1/2 [40]. This corresponds to a global (modified or not) irregular class of the form
$\bm{\Theta}^{III(0)}_{std}=\bm{\breve{\Theta}}^{III(0)}_{std}=\langle\alpha z^{1/2}\rangle_{\infty}+\langle\beta z^{-1/2}\rangle_{a},$
with $a\neq\infty$, $\alpha\neq 0$, $\beta\neq 0$.
The corresponding forest is
@@ 4.5. Doubly degenerate Painlevé III
Let us study the orbit $\operatorname{SL}_{2}(\mathbb{C})\cdot\bm{\breve{\Theta}}^{III(0)}_{std}$. The generic representation has rank 5 and admits representatives of the form
$\bm{\breve{\Theta}}^{III(0)}_{gen}=\langle\lambda z^{2}+\alpha z^{1/3}\rangle_{\infty}+\langle\mu z^{2}+\beta z^{1/2}\rangle_{\infty},$
with $\lambda\neq\mu$, $\alpha\neq 0$, $\beta\neq 0$.
The generic tree is
111122221/2121/21 / 21/3131/31 / 3
We have $k=2$ Fourier sphere coefficients, $N_{1}=N_{1}^{-}=\{\langle\lambda z^{2}+\alpha z^{1/3}\rangle_{\infty}\}$ and $N_{2}=N_{2}^{-}=\{\langle\mu z^{2}+\beta z^{1/2}\rangle_{\infty}\}$. Since the principal subtrees are not isomorphic, the two nongeneric forests are distinct. Besides the class of the standard representation, corresponding to $i=1$, the other nongeneric one has rank 3 and admits representatives of the form
$\bm{\breve{\Theta}}^{III(0)}_{2}=\langle\alpha z^{1/3}\rangle_{\infty}+\langle\beta z^{-1}\rangle_{0},$
with $\alpha\neq 0$, $\beta\neq 0$.
Finally the diagram is
666611-1- 133-3- 3
and the three classes of representations correspond to different readings of the diagram, as indicated on the figure below:
@@ 4.6. Painlevé IV
The standard rank 2 Painlevé IV Lax representation has an irregular singularity at infinity corresponding to a pole of order 3, and a regular singularity at finite distance.
The corresponding fission forest is
11112222
Any connection of this type can be brought, by applying a twist to kill a part of the formal monodromy at the regular singularity, to a connection with a modified irregular class of the form
@@
$\bm{\breve{\Theta}}^{IV}_{std}=\langle\alpha z^{2}\rangle_{\infty}+\langle\beta z^{2}\rangle_{\infty}+\langle 0\rangle_{a},$
with $\alpha\neq\beta$.
Let us describe the classes of representations of $\bm{\breve{\Theta}}^{IV}_{std}$. The generic representation has rank 3. It admits representatives of the form
$\bm{\breve{\Theta}}^{IV}_{gen}=\langle\alpha z^{2}\rangle_{\infty}+\langle\beta z^{2}\rangle_{\infty}+\langle\gamma z^{2}\rangle_{\infty},$
with $\alpha,\beta,\gamma$ pairwise distinct. It coincides with the rank 3 Lax representation found in [34].
The fission tree of the generic representation is
22221111
We have $k=3$ Fourier sphere coefficients, $N_{1}=N_{1}^{-}=\{\langle\alpha z^{2}\rangle_{\infty}\}$, $N_{2}=N_{2}^{-}=\{\langle\beta z^{2}\rangle_{\infty}\}$, and $N_{3}=N_{3}^{-}=\{\langle\gamma z^{2}\rangle_{\infty}\}$.
We observe that the three principal subtrees of the generic fission forest are isomorphic. This implies that the forests of the three nongeneric representations coincide (this reflects the fact that $\alpha,\beta,\gamma$ have entirely symmetric roles), so there are only two distinct forests, the generic one and a nongeneric one.
Finally, the diagram is the following:
and the two forests correspond to different types of readings of the diagram, as indicated on the figure below:
$\langle\alpha z^{2}\rangle_{\infty}$$\langle\beta z^{2}\rangle_{\infty}$$\langle\gamma z^{2}\rangle_{\infty}$$\bm{\breve{\Theta}}^{IV}_{gen}$$\langle\alpha z^{2}\rangle_{\infty}$$\langle\beta z^{2}\rangle_{\infty}$$\langle 0\rangle_{a}$$\bm{\breve{\Theta}}^{IV}_{std}$
4.7. Painlevé V
The standard Painlevé V Lax representation has rank 2, one irregular singularity at infinity corresponding to a pole of order 2, and two regular singularities at finite distance.
The corresponding fission forest is
$1$
If $(E,\nabla)$ is a connection of this type, we can always apply a twist to kill a part of the formal monodromy at the regular singularities, to obtain a connection with a modified global irregular class of the form:
$\bm{\breve{\Theta}}^{V}_{std}=\langle\alpha z\rangle_{\infty}+\langle\beta z\rangle_{\infty}+\langle 0\rangle_{a}+\langle 0\rangle_{b},$
with $\alpha\neq\beta$, $a\neq b$.
Let us describe the classes of representations of $\bm{\breve{\Theta}}^{V}_{std}$. The generic representation has rank 4 and corresponds to irregular classes of the form
$\bm{\breve{\Theta}}^{V}_{gen}=\langle\lambda z^{2}+\alpha z\rangle_{\infty}+\langle\lambda z^{2}+\beta z\rangle_{\infty}+\langle\mu z^{2}+az\rangle_{\infty}+\langle\mu z^{2}+bz\rangle_{\infty}$
with $\lambda\neq\mu$, $\alpha\neq\beta$, $a\neq b$.
The generic fission tree is thus
11112222
We have $k=2$ Fourier sphere coefficients, $N_{1}=N_{1}^{-}=\{\langle\lambda z^{2}+\alpha z\rangle_{\infty},\langle\lambda z^{2}+\beta z\rangle_{\infty}\}$ and $N_{2}=N_{2}^{-}=\{\langle\mu z^{2}+az\rangle_{\infty},\langle\mu z^{2}+bz\rangle_{\infty}\}$. Furthermore, the two principal subtrees corresponding to $N_{1}$ and $N_{2}$ are isomorphic, so the two nongeneric classes of representations have in fact the same forest. In turn, there are only two distinct forests of basic representations, the generic forest and the nongeneric one.
The diagram associated to $\bm{\breve{\Theta}}^{V}_{std}$ is
and the two forests correspond to different types of readings of the diagram, as indicated on the figure below:
$\langle\lambda z^{2}+\alpha z\rangle_{\infty}$$\langle\lambda z^{2}+\beta z\rangle_{\infty}$$\langle\mu z^{2}+az\rangle_{\infty}$$\langle\mu z^{2}+bz\rangle_{\infty}$$\bm{\breve{\Theta}}^{V}_{gen}$$\langle\alpha z\rangle_{\infty}$$\langle\beta z\rangle_{\infty}$$\langle 0\rangle_{a}$$\langle 0\rangle_{b}$$\bm{\breve{\Theta}}^{V}_{std}$
4.8. Painlevé VI
The standard rank 2 Lax representation for Painlevé VI has 4 regular singularities, one of them at infinity. It corresponds to a global irregular class of the form
$\bm{\Theta}^{VI}_{std}=2\langle 0\rangle_{\infty}+2\langle 0\rangle_{a}+2\langle 0\rangle_{b}+2\langle 0\rangle_{c}.$
with poles at finite distance $a,b,c,$ pairwise distinct.
The fission forest is
Furthermore, the formal monodromy at every singularity has two distinct eigenvalues. Up to tensoring with a rank one local system with regular singularities at $a,b,c$, we may assume that one of these eigenvalues is equal to 1. The corresponding modified irregular class is then
$\bm{\breve{\Theta}}^{VI}_{std}=2\langle 0\rangle_{\infty}+\langle 0\rangle_{a}+\langle 0\rangle_{b}+\langle 0\rangle_{c}.$
Let us describe the classes of nearby representations of $\bm{\breve{\Theta}}^{VI}_{std}$. The generic class of representations can be reached from $\bm{\breve{\Theta}}^{VI}_{std}$ by applying a twist followed by the Fourier transform. It has rank 5 and corresponds to irregular classes of the form:
@@ 4.8. Painlevé VI
$\bm{\breve{\Theta}}^{VI}_{gen}=2\langle\lambda z^{2}\rangle_{\infty}+\langle\mu z^{2}+\alpha z\rangle_{\infty}+\langle\mu z^{2}+\beta z\rangle_{\infty}+\langle\mu z^{2}+\gamma z\rangle_{\infty},$
with $\lambda\neq\mu$, $\alpha,\beta,\gamma$ pairwise distinct.
The corresponding fission tree is
111122222222
The number of Fourier sphere coefficients is $k=2$, and we have $N_{1}=N_{1}^{-}=\{\langle\lambda z^{2}\rangle\}$, $N_{2}=N_{2}^{-}=\{\langle\mu z^{2}+\beta z\rangle,\langle\mu z^{2}+\gamma z\rangle\}$.
In turn there are three distinct classes of representations. The standard representation corresponds to the nongeneric class of representations for $i=1$. The other nongeneric class of representations, for $i=2$, has rank 3 and admits representatives of the form
$\bm{\breve{\Theta}}^{VI}_{2}=2\langle 0\rangle_{0}+\langle\alpha z\rangle_{\infty}+\langle\beta z\rangle_{\infty}+\langle\gamma z\rangle_{\infty},$
with $\alpha,\beta,\gamma$ pairwise distinct. It corresponds to the Harnad dual [28] of the standard Lax representation.
The diagram associated to $\bm{\breve{\Theta}}^{VI}_{std}$ is the affine D4 Dynkin diagram drawn below
where the right node is not coloured to indicate that it does not belong to the core diagram, but comes from the conjugacy class of the formal monodromy of the Stokes circle with multiplicity 2, corresponding to the central vertex.
The classes of representations correspond to three different readings of the diagram, as indicated on the following figure:
@@
$\langle 0\rangle_{a}$$\langle 0\rangle_{b}$$\langle 0\rangle_{c}$$\langle 0\rangle_{\infty}$$\bm{\breve{\Theta}}^{VI}_{std}$$\langle\alpha z\rangle_{\infty}$$\langle\beta z\rangle_{\infty}$$\langle\gamma z\rangle_{\infty}$$\langle 0\rangle_{0}$$\bm{\breve{\Theta}}^{VI}_{2}$$\langle\mu z^{2}+\alpha z\rangle_{\infty}$$\langle\mu z^{2}+\beta z\rangle_{\infty}$$\langle\mu z^{2}+\gamma z\rangle_{\infty}$$\langle\lambda z^{2}\rangle_{\infty}$$\bm{\breve{\Theta}}^{VI}_{gen}$

### v3

@@ 4.5.3 Diagram
4.6 Doubly degenerate Painlevé III
4.6.1 Standard representation
The standard Lax representation for the doubly degenerate Painlevé III equation has two irregular singularities, both corresponding to a second order pole, with a nilpotent leading term at each pole so that in the Turritin–Levelt decomposition, there is a single twisted Stokes circle of slope 1/2 [45].
This corresponds to the following fission forest $\mathbf{F}^{\rm III(0)}_{\rm std}$:
@@ 4.6.1 Standard representation
The irregular classes with fission forest $\mathbf{F}^{\rm III(0)}_{\rm std}$ and minimal Poincaré–Katz ranks (both equal to 1/2) are those of the form
$\bm{\Theta}^{\rm III(0)}_{\rm std}=\bigl\langle\alpha z_{a}^{-1/2}\bigr\rangle_{a}+\bigl\langle\beta z_{b}^{-1/2}\bigr\rangle_{b},$
with $\in\mathbb{P}^{1}, aba\neq b$, $\alpha\neq 0$, $\beta\neq 0$.
Given a connection of this type, the rank of its generic class of nearby representations is minimal if one of the singularities, say $\beta$, is at infinity (to which we can reduce by applying a Möbius transformation).
Let us thus consider a (modified) irregular class of the form
$\bm{\breve{\Theta}}^{\rm III(0)}_{\rm std}=\bigl\langle\alpha z_{a}^{-1/2}\bigr\rangle_{a}+\bigl\langle\beta z^{1/2}\bigr\rangle_{\infty},$
and determine its classes of nearby representations.
4.6.2 Generic representation
The short fission tree $\mathbf{T}^{\rm III(0)}$ of $\bm{\breve{\Theta}}^{\rm III(0)}_{\rm std}$ is the following:
11221/21/21/31/3
An irregular class of generic form has fission tree $\mathbf{T}^{\rm III(0)}$ if and only if it is of the form
$\bm{\breve{\Theta}}^{\rm III(0)}_{\rm gen}=\bigl\langle\lambda^{\prime}z^{2}+\alpha^{\prime}_{1}z+\alpha^{\prime}_{1/3}z^{1/3}\bigr\rangle_{\infty}+\bigl\langle\mu^{\prime}z^{2}+\beta^{\prime}_{1}z+\beta^{\prime}_{1/2}z^{1/2}\bigr\rangle_{\infty},$
with $\lambda^{\prime}\neq\mu^{\prime}$, $\alpha^{\prime}_{1/3}\neq 0$, $\beta^{\prime}_{1/2}\neq 0$.
The generic representation has thus rank 5, $k=2$ Fourier sphere coefficients, and the corresponding partition of the set of Stokes circles of $\bm{\breve{\Theta}}^{\rm III(0)}_{\rm gen}$ is given by
$N_{1}=N_{1}^{-}=\big\{\bigl\langle\lambda^{\prime}z^{2}+\alpha^{\prime}_{1}z+\alpha^{\prime}_{1/3}z^{1/3}\bigr\rangle_{\infty}\big\},\qquad N_{2}=N_{2}^{-}=\big\{\bigl\langle\mu^{\prime}z^{2}+\beta^{\prime}_{1}z+\beta^{\prime}_{1/2}z^{1/2}\bigr\rangle_{\infty}\big\}.$
4.6.3 Nongeneric representations
Since the principal subtrees are not isomorphic, the two nongeneric forests are distinct. The standard representation corresponds to the nongeneric class of representations with $i=1$.
The other nongeneric representation, for $i=2$, has the following fission forest $\mathbf{F}^{\rm III(0)}_{2}$:
1/31/3112
Explicitly, elements of the orbit $\operatorname{SL}_{2}(\mathbb{C})\cdot\bm{\breve{\Theta}}^{\rm III(0)}_{\rm gen}$ belonging to this class of representations are of the form
$\bm{\breve{\Theta}}^{\rm III(0)}_{2}=\bigl\langle\beta^{\prime\prime}z_{b^{\prime\prime}}^{-1}\bigr\rangle_{b^{\prime\prime}}+\bigl\langle\alpha^{\prime\prime}z^{1/3}\bigr\rangle_{\infty}$
with $b^{\prime\prime}\in\mathbb{P}^{1}\smallsetminus\{\infty\}$, $\alpha^{\prime\prime}\neq 0$ and $\beta^{\prime\prime}\neq 0$, hence the corresponding non-modified irregular class is
$\bm{\Theta}^{\rm III(0)}_{2}=\bigl\langle\beta^{\prime\prime}z_{b^{\prime\prime}}^{-1}\bigr\rangle_{b^{\prime\prime}}+2\bigl\langle 0\bigr\rangle_{b^{\prime\prime}}+\bigl\langle\alpha^{\prime\prime}z^{1/3}\bigr\rangle_{\infty},$
and its fission forest is indeed $\mathbf{F}^{\rm III(0)}_{2}$.
The first nongeneric class of representations has thus rank 3, and minimal Poincaré–Katz ranks $1$, $1/3$.
4.6.4 Diagram
The corresponding diagram $\Gamma^{\rm III(0)}$ is the following [20, Section 7]:
661-13-3
The three classes of representations correspond to different readings of the diagram, as indicated on the figure below
@@ 4.7.1 Standard representation
The standard rank 2 Painlevé IV Lax representation has an untwisted irregular singularity corresponding to a pole of order 3, and a regular singularity.
This corresponds to the following fission forest $\mathbf{F}^{\rm IV}_{\rm std}$:
11222
The irregular classes with fission forest $\mathbf{F}^{\rm IV}_{\rm std}$ and minimal Poincaré–Katz ranks are those of the form
$\bm{{\Theta}}^{\rm IV}_{\rm std}=2\bigl\langle 0\bigr\rangle_{a}+\bigl\langle\alpha_{2}z_{b}^{-2}+\alpha_{1}z_{b}^{-1}\bigr\rangle_{b}+\bigl\langle\beta_{2}z_{b}^{-2}+\beta_{1}z_{b}^{-1}\bigr\rangle_{b}$
with $\in\mathbb{P}^{1}, aba\neq b$, $\alpha_{2},\alpha_{1},\beta_{2},\beta_{1}\in\mathbb{C}$, and $\alpha_{2}\neq\beta_{2}$.
Given a connection of this type, its generic class of nearby representations is of minimal rank if $b=\infty$, and its formal monodromy for $\bigl\langle 0\bigr\rangle_{b}$ admits 1 as an eigenvalue (we can always reduce to this situation by applying a Möbius transformation and a twist). This corresponds to having a modified irregular class of the form
@@
$\bm{\breve{\Theta}}^{\rm IV}_{\rm std}=\bigl\langle 0\bigr\rangle_{a}+\bigl\langle\alpha_{2}z^{2}+\alpha_{1}z\bigr\rangle_{\infty}+\bigl\langle\beta_{2}z^{2}+\beta_{1}z\bigr\rangle_{\infty}$
with $a\in\mathbb{P}^{1}\smallsetminus\{\infty\}$, and $\alpha_{2}\neq\beta_{2}$. Let us thus determine the classes of nearby representations of $\bm{\breve{\Theta}}^{\rm IV}_{\rm std}$.
4.7.2 Generic representation
The short fission tree $\mathbf{T}^{\rm IV}$ of $\bm{\breve{\Theta}}^{\rm IV}_{\rm std}$ is the following:
2211
An irregular class of generic form has fission tree $\mathbf{T}^{\rm IV}$ if and only if it is of the form
@@ 4.7.2 Generic representation
$\bm{\breve{\Theta}}^{\rm IV}_{\rm gen}=\bigl\langle\lambda^{\prime}z^{2}+\alpha^{\prime}z\bigr\rangle_{\infty}+\bigl\langle\mu^{\prime}z^{2}+\beta^{\prime}z\bigr\rangle_{\infty}+\bigl\langle\nu^{\prime}z^{2}+\gamma^{\prime}z\bigr\rangle_{\infty},$
with $\lambda^{\prime}$, $\mu^{\prime}$, $\nu^{\prime}$ pairwise distinct.
Hence the generic representation has rank $3$, $k=3$ Fourier sphere coefficients, and the corresponding partition of the set of Stokes circles of $\bm{\breve{\Theta}}^{\rm IV}_{\rm gen}$ is given by
@@
$N_{1}=N_{1}^{-}=\bigl\langle\lambda^{\prime}z^{2}+\alpha^{\prime}z\bigr\rangle_{\infty},\qquad\!\!\!N_{2}=N_{2}^{-}=\bigl\langle\mu^{\prime}z^{2}+\beta^{\prime}z\bigr\rangle_{\infty},\!\!\!\qquad N_{3}=N_{3}^{-}=\bigl\langle\nu^{\prime}z^{2}+\gamma^{\prime}z\bigr\rangle_{\infty}.$
Notice that the generic class of representations corresponds to the rank 3 Lax representation for Painlevé IV found in [36].
We observe that the three principal subtrees of the generic fission forest are isomorphic. This implies that the forests of the three nongeneric representations coincide so there are only two distinct forests, the generic one and a nongeneric one, corresponding to the standard representation.
4.7.3 Diagram
The corresponding diagram $\Gamma^{\rm IV}$ is the affine $A_{2}$ Dynkin diagram [11, 20]
The two classes of representations correspond to different readings of the diagram, as indicated on the figure below
$\bigl\langle\lambda^{\prime}z^{2}+\alpha^{\prime}z\bigr\rangle_{\infty}$$\bigl\langle\mu^{\prime}z^{2}+\beta^{\prime}z\bigr\rangle_{\infty}$$\bigl\langle\nu^{\prime}z^{2}+\gamma^{\prime}z\bigr\rangle_{\infty}$$\bm{\breve{\Theta}}^{\rm IV}_{\rm gen}$$\bigl\langle\alpha_{2}z^{2}+\alpha_{1}z\bigr\rangle_{\infty}$$\bigl\langle\beta_{2}z^{2}+\beta_{1}z\bigr\rangle_{\infty}$$\bigl\langle 0\bigr\rangle_{a}$$\bm{\breve{\Theta}}^{\rm IV}_{\rm std}$
4.8 Painlevé V
4.8.1 Standard representation
The standard Painlevé V Lax representation has rank 2, one irregular singularity corresponding to a pole of order 2, and two regular singularities.
The corresponding fission forest $\mathbf{F}^{\rm V}_{\rm std}$ is the following:
1122
The irregular classes with fission forest $\mathbf{F}^{\rm V}_{\rm std}$ and minimal Poincaré–Katz rank are exactly those of the form
$\bm{{\Theta}}^{\rm V}_{\rm std}=2\bigl\langle 0\bigr\rangle_{a}+2\bigl\langle 0\bigr\rangle_{b}+\bigl\langle\alpha z_{c}^{-1}\bigr\rangle_{c}+\bigl\langle\beta z_{c}^{-1}\bigr\rangle_{c}$
with $a,b,c\in\mathbb{P}^{1}$ pairwise distinct, $\alpha,\beta\in\mathbb{C}$, $\alpha\neq\beta$.
If $(E,\nabla)$ is a connection of this type, its generic class of nearby representations is of minimal rank if $c=\infty$, and if the conjugacy class in $\operatorname{GL}_{2}(\mathbb{C})$ of its formal monodromy for the tame circles $\bigl\langle 0\bigr\rangle_{a}$ and $\bigl\langle 0\bigr\rangle_{b}$ has $1$ as an eigenvalue (we can always reduce to this case by applying a Möbius transformation and a twist). This corresponds to having a modified irregular class of the form
$\bm{\breve{\Theta}}^{\rm V}_{\rm std}=\bigl\langle 0\bigr\rangle_{a}+\bigl\langle 0\bigr\rangle_{b}+\bigl\langle\alpha z\bigr\rangle_{\infty}+\bigl\langle\beta z\bigr\rangle_{\infty},$
with $\alpha\neq\beta$. Let us thus describe the classes of nearby representations of $\bm{\breve{\Theta}}^{\rm V}_{\rm std}$.
4.8.2 Generic representation
The short fission tree $\mathbf{T}^{\rm V}$ of $\bm{\breve{\Theta}}^{\rm V}_{\rm std}$ is
1122
An irregular class of generic form has fission tree $\mathbf{T}^{\rm V}$ if and only if it is of the form
$\bm{\breve{\Theta}}^{\rm V}_{\rm gen}=\bigl\langle\lambda^{\prime}z^{2}+\alpha^{\prime}z\bigr\rangle_{\infty}+\bigl\langle\lambda^{\prime}z^{2}+\beta^{\prime}z\bigr\rangle_{\infty}+\bigl\langle\mu^{\prime}z^{2}+\gamma^{\prime}z\bigr\rangle_{\infty}+\bigl\langle\mu^{\prime}z^{2}+\delta^{\prime}z\bigr\rangle_{\infty}$
with $\lambda^{\prime}\neq\mu^{\prime}$, $\alpha^{\prime}\neq\beta^{\prime}$, $\gamma^{\prime}\neq\delta^{\prime}$.
The generic class of representations has thus rank 4, $k=2$ Fourier sphere coefficients, and the corresponding partition of the set of Stokes circles of $\bm{\breve{\Theta}}^{\rm V}_{\rm gen}$ is given by
$N_{1}=N_{1}^{-}=\big\{\bigl\langle\lambda^{\prime}z^{2}+\alpha^{\prime}z\bigr\rangle_{\infty},\bigl\langle\lambda^{\prime}z^{2}+\beta^{\prime}z\bigr\rangle_{\infty}\big\},$
$N_{2}=N_{2}^{-}=\big\{\bigl\langle\mu^{\prime}z^{2}+\gamma^{\prime}z\bigr\rangle_{\infty},\bigl\langle\mu^{\prime}z^{2}+\delta^{\prime}z\bigr\rangle_{\infty}\big\}.$
Furthermore, the two principal subtrees corresponding to $N_{1}$ and $N_{2}$ are isomorphic, so the two nongeneric classes of representations have in fact the same forest. In turn, there are only two distinct classes of basic representations, the generic one and the nongeneric one, corresponding to the standard Lax representation.
4.8.3 Diagram
The corresponding diagram $\Gamma^{V}$ is the affine $A_{3}$ Dynkin diagram [11, 20]
The two classes of representations correspond to different readings of the diagram, as indicated on the figure below
$\bigl\langle\lambda^{\prime}z^{2}+\alpha^{\prime}z\bigr\rangle_{\infty}$$\bigl\langle\lambda^{\prime}z^{2}+\beta^{\prime}z\bigr\rangle_{\infty}$$\bigl\langle\mu^{\prime}z^{2}+\gamma^{\prime}z\bigr\rangle_{\infty}$$\bigl\langle\mu^{\prime}z^{2}+\delta^{\prime}z\bigr\rangle_{\infty}$$\bm{\breve{\Theta}}^{\rm V}_{\rm gen}$$\bigl\langle\alpha z\bigr\rangle_{\infty}$$\bigl\langle\beta z\bigr\rangle_{\infty}$$\bigl\langle 0\bigr\rangle_{a}$$\bigl\langle 0\bigr\rangle_{b}$$\bm{\breve{\Theta}}^{\rm V}_{\rm std}$
4.9 Painlevé VI
4.9.1 Standard representation
The standard rank 2 Lax representation for Painlevé VI has 4 regular singularities.
The corresponding fission forest $\mathbf{F}^{\rm VI}_{\rm std}$ is the following (all fission trees are trivial):
$22$
The irregular classes with fission forest $\mathbf{F}^{\rm VI}_{\rm std}$ are those of the form
$\bm{\Theta}^{\rm VI}_{\rm std}=2\bigl\langle 0\bigr\rangle_{a}+2\bigl\langle 0\bigr\rangle_{b}+2\bigl\langle 0\bigr\rangle_{c}+2\bigl\langle 0\bigr\rangle_{d}.$
with $a,b,c,d\in\mathbb{P}^{1}$ pairwise distinct.
Given a connection of this type, its generic class of nearby representations has minimal rank if one of the singularities, say $d$, is at infinity, and if for the other singularities $a$, $b$, $c$, the formal monodromy of the connection has 1 as an eigenvalue (once again we can always reduce to this situation by applying a Möbius transformation and a twist). This corresponds to a modified irregular class of the form
$\bm{\breve{\Theta}}^{\rm VI}_{\rm std}=2\bigl\langle 0\bigr\rangle_{\infty}+\bigl\langle 0\bigr\rangle_{a}+\bigl\langle 0\bigr\rangle_{b}+\bigl\langle 0\bigr\rangle_{c},$
with $a,b,c\in\mathbb{P}^{1}\smallsetminus\{\infty\}$ pairwise distinct. Let us thus determine the classes of nearby representations of $\bm{\breve{\Theta}}^{\rm VI}_{\rm std}$.
4.9.2 Generic representation
The short fission tree $\mathbf{T}^{\rm VI}$ of $\bm{\breve{\Theta}}^{\rm VI}_{\rm std}$ is
112222
An irregular class of generic form has fission tree $\mathbf{T}^{\rm VI}$ if and only if it is of the form
$\bm{\breve{\Theta}}^{\rm VI}_{\rm gen}=2\bigl\langle\lambda^{\prime}z^{2}+\alpha^{\prime}z\bigr\rangle_{\infty}+\bigl\langle\mu^{\prime}z^{2}+\beta^{\prime}z\bigr\rangle_{\infty}+\bigl\langle\mu^{\prime}z^{2}+\gamma^{\prime}z\bigr\rangle_{\infty}+\bigl\langle\mu^{\prime}z^{2}+\delta^{\prime}z\bigr\rangle_{\infty},$
with $\lambda^{\prime}\neq\mu^{\prime}$, and $\beta^{\prime}$, $\gamma^{\prime}$, $\delta^{\prime}$ pairwise distinct.
The generic class of representations has rank $5$, its number of Fourier sphere coefficients is k=2{k=2}, and the corresponding partition of the set of Stokes circles of $\bm{\breve{\Theta}}^{\rm VI}_{\rm gen}$ is given by
$N_{1}=N_{1}^{-}=\big\{\bigl\langle\lambda^{\prime}z^{2}+\alpha^{\prime}z\bigr\rangle\big\},$
$N_{2}=N_{2}^{-}=\big\{\bigl\langle\mu^{\prime}z^{2}+\beta^{\prime}z\bigr\rangle,\bigl\langle\mu^{\prime}z^{2}+\gamma^{\prime}z\bigr\rangle,\bigl\langle\mu^{\prime}z^{2}+\delta^{\prime}z\bigr\rangle\big\}.$
4.9.3 Nongeneric representations
Since the two principal subtrees are not isomorphic, there are two distinct classes of nongeneric nearby representations. The standard representation corresponds to the nongeneric class of representations for $i=2$.
The other nongeneric class of representations, for $i=1$, has the following fission forest $\mathbf{F}^{\rm VI}_{1}$:
112233
Explicitly, elements of the orbit $\operatorname{SL}_{2}(\mathbb{C})\cdot\bm{\breve{\Theta}}^{\rm VI}_{\rm gen}$ belonging to this class of representations are of the form
$\bm{\breve{\Theta}}^{\rm VI}_{1}=2\bigl\langle 0\bigr\rangle_{a^{\prime\prime}}+\bigl\langle\alpha^{\prime\prime}z\bigr\rangle_{\infty}+\bigl\langle\beta^{\prime\prime}z\bigr\rangle_{\infty}+\bigl\langle\gamma^{\prime\prime}z\bigr\rangle_{\infty},$
with $a^{\prime\prime}\in\mathbb{P}^{1}\smallsetminus\{\infty\}$, and $\alpha^{\prime\prime},\beta^{\prime\prime},\gamma^{\prime\prime}\in\mathbb{C}$ pairwise distinct. The corresponding non-modified irregular class is thus
$\bm{\Theta}^{\rm VI}_{1}=3\bigl\langle 0\bigr\rangle_{a^{\prime\prime}}+\bigl\langle\alpha^{\prime\prime}z\bigr\rangle_{\infty}+\bigl\langle\beta^{\prime\prime}z\bigr\rangle_{\infty}+\bigl\langle\gamma^{\prime\prime}z\bigr\rangle_{\infty},$
and its fission forest is indeed $\mathbf{F}^{\rm VI}_{1}$.
Notice that this class of representations corresponds to the well-known Harnad dual [30] of the standard Lax representation.
4.9.4 Diagram
The corresponding diagram $\Gamma^{\rm VI}$ is the affine $D_{4}$ Dynkin diagram famously associated to Painlevé VI
Here, the right node is not coloured, to indicate that it does not belong to the core diagram, but comes from the conjugacy class of the formal monodromy of the Stokes circle with multiplicity 2, corresponding to the central vertex.
The classes of representations correspond to three different readings of the diagram, as indicated on the figure below
$\bigl\langle 0\bigr\rangle_{a}$$\bigl\langle 0\bigr\rangle_{b}$$\bigl\langle 0\bigr\rangle_{c}$$\bigl\langle 0\bigr\rangle_{\infty}$$\bm{\breve{\Theta}}^{\rm VI}_{\rm std}$$\bigl\langle\alpha^{\prime\prime}z\bigr\rangle_{\infty}$$\bigl\langle\beta^{\prime\prime}z\bigr\rangle_{\infty}$$\bigl\langle\gamma^{\prime\prime}z\bigr\rangle_{\infty}$$\bigl\langle 0\bigr\rangle_{a^{\prime\prime}}$$\bm{\breve{\Theta}}^{\rm VI}_{1}$$\bigl\langle\mu^{\prime}z^{2}+\beta^{\prime}z\bigr\rangle_{\infty}$$\bigl\langle\mu^{\prime}z^{2}+\gamma^{\prime}z\bigr\rangle_{\infty}$$\bigl\langle\mu^{\prime}z^{2}+\delta^{\prime}z\bigr\rangle_{\infty}$$\bigl\langle\lambda^{\prime}z^{2}+\alpha^{\prime}z\bigr\rangle_{\infty}$$\bm{\breve{\Theta}}^{\rm VI}_{\rm gen}$